# Closed Orbits of Dynamically Convex Reeb Flows: Towards the HZ- and Multiplicity Conjectures

### v2

@@ 1.1. Introduction
In various settings the conjecture is far from new. For instance, Part (C-M), often referred to as the multiplicity problem, is stated in, say, [ALM] in full generality exactly as above. For convex hypersurfaces in ${\mathbb{R}}^{2n}$, it is included in Ekeland’s monograph, [Ek, p. 198], and apparently it was a well-known open question already then. Thus, since convex hypersurfaces are dynamically convex (see [HWZ]), our first result establishes Ekeland’s variant of the conjecture and, in fact, is more general. (Dynamical convexity is the lower bound $n+1$, the same as for ellipsoids, for the Conley–Zehnder index of closed orbits. This condition is in general less restrictive than convexity; see [CE22a, CE22b, CGHi, DGRZ] and also [AM22, GiM].) For non-degenerate flows, Ekeland’s conjecture was proved in [LZ] (see also [Lo02]) for convex hypersurfaces and in [GK] in the dynamically convex case. Recently, again in the non-degenerate setting, the dynamical convexity condition has been significantly relaxed to a very minor restriction on the index; see [DL2W18, DL2W22, GGM] and references therein.
Without the non-degeneracy requirement, the previously known lower bound on the number of prime closed orbits is only $\lceil n/2\rceil+1$, i.e., roughly $n/2$, and the full form of dynamical convexity is used in the proofs; see [DL, GG20] and also [LZ, Wa16b] for earlier results. As we have already mentioned, in Theorem A we improve this lower bound to $n$, i.e., by roughly a factor of two. (When $n=3$ or $4$, the lower bound $3$ has been established in [Wa16a, WHL].)
@@
Finally, the case of dimension $32n-1=$ 3 deserves a separate discussion. According to the so called 2-or-infinity conjecture, a Reeb flow on a closed 3-manifold has either exactly two or infinitely many prime closed orbits. This conjecture is almost proved in general, and lens spaces are the only manifolds admitting Reeb pseudo-rotations; see [CGHu, CGHP, CGH2L, CDR]. Furthermore, the conjecture has been completely proved for $S^{3}$. (See also [GH2M] for a different proof of (C-M) for $S^{3}$.) Hence, both (C-M) and (C-HZ) parts of our $n$-or-$\infty$ conjecture hold in this case. However, one cannot expect this clear-cut dichotomy to hold in higher dimensions, even for tight or fillable contact manifolds. Indeed, while it is reasonable to expect that a majority of such contact manifolds do not admit Reeb pseudo-rotations in all dimensions just as in dimension three, it is not clear how to rigorously state this conjecture and a description of manifolds that do seems beyond our reach. Moreover, the lower bound on the number of orbits clearly depends on the manifold when $n>2$.
@@
The key to the proofs of the two main theorems is Theorem C asserting that for a dynamically convex Reeb pseudo-rotation the (non-equivariant) filtered symplectic homology is one-dimensional for every action threshold and any ground field. The proof of this theorem has three main ingredients. The first one is a new version of the index recurrence theorem (Theorem 3.3), refining the original index recurrence theorem from [GG20, Thm. 5.2] which, in turn, is closely related to the common jump theorem from [Lo02, LZ]. This is a combinatorial result about, as the name suggests, recurrent behavior of the Conley–Zehnder index type invariants under iterations. This theorem and its consequences are repeatedly used throughout the paper. The other two ingredients are not new: one is an a priori upper bound on the boundary depth of the symplectic homology persistence module (Theorem 2.3, originally from [GS]) and the other is the Smith inequality for symplectic homology (Theorem 2.5, originally from [Se15, ShZh]). Here, due to the nature of the Smith inequality which is used to obtain a lower bound on the dimension, we first need to work over fields of positive characteristic and then extend the result to all fields by the universal coefficient theorem. Once Theorem C is proved, we can set the ground field to be ${\mathbb{Q}}$ or any field of zero characteristic.
@@ 1.2. Main results
When $x$ is non-degenerate, we call it non-alternating if the number, counted with algebraic multiplicity, of real negative eigenvalues in $(-1,0)$ of its Poincaré return map is even. Otherwise, we call $x$ alternating. A non-degenerate even iterate of an alternating closed orbit is called bad; all other non-degenerate closed orbits are called good. When $M$ is centrally symmetric, we call $x$ symmetric if $-x(t)=x(t+T/2)$, where $T$ is the period of $x$, i.e., $x$ and $-x$ are the same orbit up to the choice of initial condition.
@@
This theorem is proved in Section 4.1. The key new point here is, of course, (M1), establishing Ekeland’s conjecture and improving previously known multiplicity lower bounds by roughly a factor of two. As we have already mentioned, the existence of at least $n$ non-alternating orbits for non-degenerate dynamically convex Reeb flows is well-known; see [GK, LZ]. The new part in (M2), essential for the proof of Theorem B, is that in this case the number of non-alternating closed prime orbits is exactly $n$ and the remaining prime orbits, if they exist, must all be alternating.
@@
Turning to our more technical results, fix a background field ${\mathbb{F}}$. When ${\mathbb{F}}$ has positive characteristic, ${\mathit{c}har}\,{\mathbb{F}}>0$, the reader may take ${\mathbb{F}}={\mathbb{Z}}_{p}$ and when ${\mathit{c}har}\,{\mathbb{F}}=0$ it would be sufficient, without loss of generality, to set ${\mathbb{F}}={\mathbb{Q}}$. Denote by $\operatorname{SH}^{t}(W;{\mathbb{F}})$, where the $t\in{\mathbb{R}}$ is the action or equivalently period threshold, the filtered symplectic homology of $W$ with coefficients in ${\mathbb{F}}$. For each $t$, this is a ${\mathbb{Z}}$-graded vector space over ${\mathbb{F}}$. Throughout the paper it is convenient to treat the family of vector spaces $\operatorname{SH}^{t}(W;{\mathbb{F}})$ as a graded persistence module, which we denote by $\operatorname{SH}(W;{\mathbb{F}})$; see Sections 2.2 and 2.3. The next result, proved in Section 4.3, is central to the proofs of Theorems A and B.
@@
From now on, we can restrict our attention to ground fields of zero characteristic, e.g., ${\mathbb{F}}={\mathbb{Q}}$. In what follows, we call $h$ from (1.2) the degree $\deg(x)$ of $x$. In other words, $h=\deg(x)\in{\mathbb{Z}}$ is the only integer such that $\operatorname{CH}_{h}(x;{\mathbb{Q}})\neq 0$ (and hence is ${\mathbb{Q}}$.) For instance, $\deg(x)=\mu(x)$ when $x$ is non-degenerate and good. We emphasize that this notion is specific to the setting of Theorem C and Corollary D, i.e., $\deg(x)$ is defined only when the flow is dynamically convex and has finitely many prime closed Reeb orbits and $x$ is ${\mathbb{Q}}$-visible. Recall that ${\mathcal{P}}={\mathcal{P}}({\mathbb{Q}})$ stands for the set of all, not necessarily prime, ${\mathbb{Q}}$-visible closed orbits. Denote by ${\mathcal{A}}(x)$ the action (aka the period) of $x$ and by $\operatorname{\hat{\mu}}(x)$ the mean index of $x$; see Sections 2.1 and 3.1.
@@
This theorem asserts, in particular, that the graded barcode of $\operatorname{SH}(W;{\mathbb{Q}})$ looks exactly like that of an irrational ellipsoid: the bars form an ascending infinite staircase with one riser of height one (the degree) for each closed orbit and starting with a tread of degree $n$ beginning at the interior of $W$. We will show in the next section that the similarity with ellipsoids goes farther. Note also that as a consequence of (O1) the flow is automatically lacunary in the terminology from [AM24].
@@ 1.3. Application: ellipsoid comparison
Let $W\subset{\mathbb{R}}^{2n}$ be a star-shaped domain. Assume that the flow on $M=\partial W$ is non-degenerate, dynamically convex, has exactly $n$ prime closed orbits and these orbits are non-alternating. In addition, let us require the following non-resonance condition: for every $j$ all ratios ±$\pm\Delta_{j}/\Delta_{i}$, $i\neq j$, are distinct modulo ${\mathbb{Z}}$. Then the linearized flows along the corresponding closed orbits in $M$ and $M_{\vec{\Delta}}$ are equal up to conjugation as elements of $\operatorname{\widetilde{Sp}}\big{(}2(n-1)\big{)}$.
@@ Acknowledgments
The authors would like to thank Leonardo Macarini and Marco Mazzucchelli for useful remarks and suggestions.
@@ 2.1. Conventions and notation
[…]
@@
Recall also that a non-degenerate closed orbit $x$ is said to be non-alternating if the number (counted with algebraic multiplicity) of eigenvalues in $(-1,0)$ of the return map is even. Otherwise, we call $x$ alternating. When $x$ is strongly non-degenerate, these conditions are equivalent to that all indices $\mu\big{(}x^{k}\big{)}$, $k\in{\mathbb{N}}$, have the same parity or that the parity alternates. (A word is due on terminology: Non-alternating orbits are sometimes referred to as negative hyperbolic, but we find this term misleading when $\dim M>3$. The frequently used terms “even/odd” are more intuitive; however they clash with the parity of an orbit.) While formally speaking these definitions require no conditions on $x$, we will apply them only when $x$ is prime and strongly non-degenerate. The notion of (non-)alternating closed orbit is not particularly useful when $x^{k}$ is allowed to be degenerate; see Example 2.13. The parity of a closed non-degenerate orbit $x$ is simply the parity of $\mu(x)$. A non-degenerate closed orbit $x$ is said to be bad if it is an even iterate of an alternating prime orbit, and good otherwise.
@@
This condition is obviously satisfied in two cases we are interested in here: when $\alpha$ is non-degenerate and when the Reeb flow has only finitely many prime closed orbits. The role of this condition is rather expository than conceptual, and most of the discussion in the next two sections carries over to the general case with only technical changes. For instance, the symplectic homology is still a persistence module (see Section 2.2) when ${\mathcal{S}}$ is only closed and nowhere dense and $V^{s}$ might be infinite-dimensional, once (PM2) in the definition of a persistence module is suitably modified; [ÇG2M24].
We implicitly adopt the conventions from [ÇG2M23, ÇG2M24, GG20] on the almost complex structures, Hamiltonians and the Floer equation. For the sake of brevity, since these conventions are never explicitly used in the paper, we leave them out. Here we only mention that with these conventions, the Conley–Zehnder index of a non-degenerate minimum with small Hessian in dimension $2n$ is $n$.
The characteristic of the ground field ${\mathbb{F}}$ is essential in some parts of the paper. When the characteristic is zero, we can and usually will take ${\mathbb{F}}={\mathbb{Q}}$. We use the notation ${\mathbb{F}}_{p}$ for the positive characteristic field ${\mathbb{Z}}/p{\mathbb{Z}}$, where $p$ is prime, or any other field of characteristic $p$.
@@ 2.2. Graded persistence modules
Recall furthermore that a (graded) interval persistence module ${\mathbb{F}}_{(a,\,b]}$, where $-\infty<a<b\leq\infty$, is defined by setting
@@
The normal form or structure theorem asserts that every persistence module meeting the above conditions can be decomposed as a direct sum of a countable collection (i.e., a countable multiset) of interval persistence modules. Moreover, this decomposition is unique up to re-ordering of the sum. We refer the reader [PRSZ] for a proof of this theorem and further references, and also, e.g., to [CZCG, ZC] for previous or related results. The multiset ${\mathcal{B}}(V)$ of intervals entering this decomposition is referred to as the barcode of $V$ and the intervals occurring in ${\mathcal{B}}(V)$ as bars. In addition, every bar is assigned a degree which is by definition the degree of the component of $V^{s}$ it comes from.
@@ 2.3. Filtered symplectic homology
The first versions of symplectic homology were introduced in [CFH, Vi99] and since then the theory has been further developed. We refer the reader to, e.g., [Abu, Bo, BO09a, BO09b, BO13, Se06] for the definitions and basic constructions, which we mainly omit or just informally sketch for the sake of brevity, focusing on lesser known facts. Here, treating equivariant and non-equivariant symplectic homology as persistence modules, we closely follow [ÇG2M24, GG20].
@@
This family of vector spaces together with the natural “inclusion” maps form a graded persistence module in the sense of Section 2.2 with ${\mathcal{S}}={\mathcal{S}}(\alpha)\cup\{0\}$. (It is essential at this point that $W$ meets the background requirement (BA). Otherwise, we would need to modify the definition of a persistence module; see [ÇG2M24].) We denote the resulting persistence module by $\operatorname{SH}(W;{\mathbb{F}})$. The global symplectic homology is then $\operatorname{SH}^{\infty}(W;{\mathbb{F}})$.
Somewhat informally, $\operatorname{SH}^{t}(W;{\mathbb{F}})$ can be defined via the following construction. As is pointed out it in Section 2.2, it suffices to define $\operatorname{SH}^{t}(W;{\mathbb{F}})$ when $t\not\in{\mathcal{S}}$. Assume first that the Reeb flow is non-degenerate. Let $H$ be an autonomous Hamiltonian on $\widehat{W}$, which is a convex, strictly monotone increasing function $h(r)$ of $r$ on $M\times[1,\infty)$ and 00 on $W$. The function $h$ is required to be linear at infinity, i.e., of the form $tr-c$ with $c>0$. Furthermore, let $\tilde{H}$ be a small Morse perturbation of $H$ supported near $W$. Then $\tilde{H}$ is a Morse–Bott Hamiltonian in the sense of [Bo, BO09a], and $\operatorname{SH}^{t}(W;{\mathbb{F}})$ is simply the Floer homology of $H$ or equivalently $\tilde{H}$; cf. [ÇG2M24, Thm. 3.5 and Cor. 3.6]. This is the homology of a complex with two generators $\hat{x}$ and $\check{x}$ of degrees, $\mu(x)+1$ and $\mu(x)$ respectively, for every closed Reeb of orbit $x$ with ${\mathcal{A}}(x)<t$ and one generator for every critical point of $\tilde{H}$; see, e.g., [Bo, BO09a, ÇG2M23]. In fact, one can trim down this complex further and make the number of generators of the latter type equal to $\dim\operatorname{H}(W;{\mathbb{F}})$. In any event, when $W$ is star-shaped we can have only one such generator, which we denote by $[W]$. When the Reeb flow is degenerate, we first take a small non-degenerate perturbation of the contact form and then apply this construction. For an interval $I=(t^{\prime},t)$ we define $\operatorname{SH}^{I}(W;\alpha)$ as the quotient of $\operatorname{SH}^{t}(W;{\mathbb{F}})$ by the image of $\operatorname{SH}^{t^{\prime}}(W;{\mathbb{F}})$.
@@
$\operatorname{SH}^{\delta}(W;{\mathbb{F}}):=\operatorname{H}(W;{\mathbb{F}})[-n].$ (2.2)
It also follows from this description that the Euler characteristic of the graded vector space $\operatorname{SH}^{t}(W;{\mathbb{F}})$ is independent of $t>0$. To be more precise, for any field ${\mathbb{F}}$ and every $t>0$, we have
$\sum_{m}(-1)^{m}\dim\operatorname{SH}^{t}_{m}(W)=(-1)^{n}\chi(W).$ (2.3)
The construction of symplectic homology carries over to ${\mathbb{Z}}$ or any other coefficient ring, and hence we have the universal coefficient theorem. In particular, for a fixed $t$, in every degree $m$ we have
@@
when the characteristic $p$ is large enough (depending on $t$).
@@
Recall in this connection that $\operatorname{SH}^{\infty}(W;{\mathbb{F}})$ vanishes, for instance, whenever $W$ is displaceable in $\widehat{W}$. For example, $\operatorname{SH}(W)^{\infty}=0$ when $W$ is a star-shaped domain in $\widehat{W}={\mathbb{R}}^{2n}$. In this case, vanishing of $\operatorname{SH}^{\infty}(W)$ is established in [Vi99]. The general case is proved in [CFO] via Rabinowitz–Floer homology and direct proofs are given in [Su16] and [GS]. Strictly speaking, Theorem 2.3 is proved in [GS] for ${\mathbb{F}}={\mathbb{Q}}$, but the argument carries over word-for-word to any unital ring.
@@
For every Liouville domain $W$ and any interval $I$,
@@
These inequalities were originally proved for the Floer homology of any Hamiltonian $H$ on $\widehat{W}$ of the form $H=tr-c$ outside a compact set with $t\not\in{\mathcal{S}}(\alpha)$, [Se15, ShZh]. The version stated here readily follows from those results combined with, for instance, [ÇG2M24, Cor. 3.6].

### v3

@@ 1.1. Introduction
In various settings the conjecture is far from new. For instance, Part (C-M), often referred to as the multiplicity problem, is stated in, say, [ALM] in full generality exactly as above. For convex hypersurfaces in ${\mathbb{R}}^{2n}$, it is included in Ekeland’s monograph, [Ek, p. 198], and apparently it was a well-known open question already then. Thus, since convex hypersurfaces are dynamically convex (see [HWZ]), our first result establishes Ekeland’s variant of the conjecture and, in fact, is more general. (Dynamical convexity is the lower bound $n+1$, the same as for ellipsoids, for the Conley–Zehnder index of closed orbits. This condition is in general less restrictive than convexity; see [CE22, CE25, CGHi, DGRZ] and also [AM22, GiM].) For non-degenerate flows, Ekeland’s conjecture was proved in [LZ] (see also [Lo02]) for convex hypersurfaces and in [GK] in the dynamically convex case. Recently, again in the non-degenerate setting, the dynamical convexity condition has been significantly relaxed to a very minor restriction on the index; see [DL2W18, DL2W24, GGM] and references therein.
Without the non-degeneracy requirement, the previously known lower bound on the number of prime closed orbits is only $\lceil n/2\rceil+1$, i.e., roughly $n/2$, and the full form of dynamical convexity is used in the proofs; see [DL, GG20] and also [LZ, Wa16b] for earlier results. As we have already mentioned, in Theorem A we improve this lower bound to $n$, i.e., by roughly a factor of two. (When $n=3$ and $4$, the lower bounds $3$ and, respectively, $4$ have been established in [Wa16a, WHL].)
@@
Finally, the case of dimension 2n1=32n-1=3 deserves a separate discussion. According to the so called 2-or-infinity conjecture, a Reeb flow on a closed 3-manifold has either exactly two or infinitely many prime closed orbits. This conjecture is almost proved in general, and lens spaces are the only manifolds admitting Reeb pseudo-rotations; see [CGHu, CGHP, CGH2L, CDR]. Furthermore, the conjecture has been completely proved for $S^{3}$. (See also [GH2M] for a different proof of (C-M) for $S^{3}$.) Hence, both (C-M) and (C-HZ) parts of our $n$-or-$\infty$ conjecture hold in this case. However, one cannot expect this clear-cut dichotomy to hold in higher dimensions, even for tight or fillable contact manifolds. Indeed, while it is reasonable to expect that a majority of such contact manifolds do not admit Reeb pseudo-rotations in all dimensions just as in dimension three, it is not clear how to rigorously state this conjecture and a description of manifolds that do seems beyond our reach. Moreover, the lower bound on the number of orbits clearly depends on the manifold when $n>2$.
@@
The key to the proofs of the two main theorems is Theorem C asserting that for a dynamically convex Reeb pseudo-rotation the (non-equivariant) filtered symplectic homology is one-dimensional for every action threshold and any ground field. The proof of this theorem has three main ingredients. The first one is a new version of the index recurrence theorem (Theorem 3.3), refining the original index recurrence theorem from [GG20, Thm. 5.2] which, in turn, is closely related to and can be derived from various variants of the common jump theorem from [DLW, Lo02, LZ]. This is a combinatorial result about, as the name suggests, recurrent behavior of the Conley–Zehnder index type invariants under iterations. This theorem and its consequences are repeatedly used throughout the paper. The other two ingredients are not new: one is an a priori upper bound on the boundary depth of the symplectic homology persistence module (Theorem 2.3, originally from [GS]) and the other is the Smith inequality for symplectic homology (Theorem 2.5, originally from [Se15, ShZh]). Here, due to the nature of the Smith inequality which is used to obtain a lower bound on the dimension, we first need to work over fields of positive characteristic and then extend the result to all fields by the universal coefficient theorem. Once Theorem C is proved, we can set the ground field to be $\mathbb{Q}$} or any field of zero characteristic.
@@ 1.2. Main results
When $x$ is non-degenerate, we call it non-alternating if the number, counted with algebraic multiplicity, of real negative eigenvalues in (1,0)(-1,0) of its Poincaré return map is even. Otherwise, we call $x$ alternating. A non-degenerate even iterate of an alternating closed orbit is called bad; all other non-degenerate closed orbits are called good. When $M$ is centrally symmetric, we call $x$ symmetric if x(t)=x(t+T/2)-x(t)=x(t+T/2), where $T$ is the period of $x$, i.e., $x$ and x-x are the same orbit up to the choice of initial condition.
@@
This theorem is proved in Section 4.1, where in fact we establish a more precise result (Theorem 4.1) identifying $n$ prime orbits meeting a certain visibility criterion. The key new point of Theorem A is, of course, (M1), establishing Ekeland’s conjecture and improving previously known multiplicity lower bounds by roughly a factor of two. As we have already mentioned, the existence of at least $n$ non-alternating orbits for non-degenerate dynamically convex Reeb flows is well-known; see [GK, LZ]. The new part in (M2), essential for the proof of Theorem B, is that in this case the number of non-alternating closed prime orbits is exactly $n$ and the remaining prime orbits, if they exist, must all be alternating.
@@
Turning to our more technical results, fix a background field $\mathbb{F}$}. When $\mathbb{F}$} has positive characteristic, ${\mathit{c}har}\,{\mathbb{F}}>0$, the reader may take ${\mathbb{F}}={\mathbb{Z}}_{p}$ and when ${\mathit{c}har}\,{\mathbb{F}}=0$ it would be sufficient, without loss of generality, to set $\mathbb{F}$$\mathbb{Q}$}. Denote by $\operatorname{SH}^{t}(W;{\mathbb{F}})$, where the $t\in{\mathbb{R}}$ is the action or equivalently period threshold, the filtered symplectic homology of $W$ with coefficients in $\mathbb{F}$}. For each $t$, this is a $\mathbb{Z}$}-graded vector space over $\mathbb{F}$}. Throughout the paper it is convenient to treat the family of vector spaces $\operatorname{SH}^{t}(W;{\mathbb{F}})$ as a graded persistence module, which we denote by $\operatorname{SH}(W;{\mathbb{F}})$; see Sections 2.2 and 2.3. The next result, proved in Section 4.3, is central to the proofs of Theorems A and B.
@@
From now on, we can restrict our attention to ground fields of zero characteristic, e.g., $\mathbb{F}$$\mathbb{Q}$}. In what follows, we call $h$ from (1.2) the degree $\deg(x)$ of $x$. In other words, $h=\deg(x)\in{\mathbb{Z}}$ is the only integer such that $\operatorname{CH}_{h}(x;{\mathbb{Q}})\neq 0$ (and hence is $\mathbb{Q}$}.) For instance, $\deg(x)=\mu(x)$ when $x$ is non-degenerate and good. We emphasize that this notion is specific to the setting of Theorem C and Corollary D, i.e., $\deg(x)$ is defined only when the flow is dynamically convex and has finitely many prime closed Reeb orbits and $x$ is $\mathbb{Q}$}-visible. Recall that ${\mathcal{P}}={\mathcal{P}}({\mathbb{Q}})$ stands for the set of all, not necessarily prime, $\mathbb{Q}$}-visible closed orbits. Denote by ${\mathcal{A}}(x)$ the action (aka the period) of $x$ and by $\operatorname{\hat{\mu}}(x)$ the mean index of $x$; see Sections 2.1 and 3.1.
@@
This theorem asserts, in particular, that the graded barcode of $\operatorname{SH}(W;{\mathbb{Q}})$ looks exactly like that of an irrational ellipsoid: the bars form an ascending infinite staircase with one riser of height one (the degree) for each closed orbit and starting with a tread of degree $n$ beginning at the interior of $W$. We will show in the next section that the similarity with ellipsoids goes farther. Note also that as a consequence of (O1) the flow is automatically lacunary in the terminology from [AM24]. It is also worth noting that the ratios ${\mathcal{A}}(x)/\operatorname{\hat{\mu}}(x)$ playing an important role in these questions were originally, to the best of our knowledge, considered in [Ek, EH].
Remark 1.3.
We conjecture that (1.1) is equivalent to that the Reeb flow is dynamically convex and has only a finite number of prime orbits at least in the non-degenerate case. It not hard to see that then (1.1) implies dynamical convexity. However, proving finiteness appears to be much more difficult. Without non-degeneracy one would also have to account for invisible orbits.
@@ 1.3. Application: ellipsoid comparison
Let $W\subset{\mathbb{R}}^{2n}$ be a star-shaped domain. Assume that the flow on $M=\partial W$ is non-degenerate, dynamically convex, has exactly $n$ prime closed orbits. In addition, let us require the following non-resonance condition: for every $j$ all ratios $\pm\Delta_{j}/\Delta_{i}$, $i\neq j$, are distinct modulo $\mathbb{Z}$}. Then the linearized flows along the corresponding closed orbits in $M$ and $M_{\vec{\Delta}}$ are equal up to conjugation as elements of $\operatorname{\widetilde{Sp}}\big(2(n-1)\big)$.
@@ Acknowledgments
The authors would like to thank Leonardo Macarini, Marco Mazzucchelli and Felix Schlenk for useful remarks and suggestions.
@@ 2.1. Conventions and notation
where $r$ is the coordinate on $1,\,\infty)$. Sometimes it is convenient to extend the function $r$ to a collar of $M=\partial W$ in $W$. Thus we can think of $\widehat{W}$ as the union of $W$ and $M\times[1-\epsilon,\,\infty)$ for small $\epsilon>0$ with $M\times[1-\epsilon,\,1]$ lying in $W$ and the symplectic form given by the same formula.
@@
Recall also that a non-degenerate closed orbit $x$ is said to be non-alternating if the number (counted with algebraic multiplicity) of eigenvalues in (1,0)(-1,0) of the return map is even. Otherwise, we call $x$ alternating. When $x$ is strongly non-degenerate, these conditions are equivalent to that all indices $\mu\big(x^{k}\big)$, $k\in{\mathbb{N}}$, have the same parity or that the parity alternates. (A word is due on terminology: Non-alternating orbits are sometimes referred to as negative hyperbolic, but we find this term misleading when $\dim M>3$. The frequently used terms “even/odd” are more intuitive; however they clash with the parity of an orbit.) While formally speaking these definitions require no conditions on $x$, we will apply them only when $x$ is prime and strongly non-degenerate. The notion of (non-)alternating closed orbit is not particularly useful when $x^{k}$ is allowed to be degenerate; see Example 2.13. The parity of a closed non-degenerate orbit $x$ is simply the parity of $\mu(x)$. A non-degenerate closed orbit $x$ is said to be bad if it is an even iterate of an alternating prime orbit, and good otherwise.
@@
This condition is obviously satisfied in two cases we are interested in here: when $\alpha$ is non-degenerate and when the Reeb flow has only finitely many prime closed orbits. The role of this condition is rather expository than conceptual, and most of the discussion in the next two sections carries over to the general case with only technical changes. For instance, the symplectic homology is still a persistence module (see Section 2.2) when $\mathcal{S}$} is only closed and nowhere dense and $V^{s}$ might be infinite-dimensional, once (PM2) in the definition of a persistence module is suitably modified, [ÇG2M25].
We implicitly adopt the conventions from [ÇG2M23, ÇG2M25, GG20] on the almost complex structures, Hamiltonians and the Floer equation. For the sake of brevity, since these conventions are never explicitly used in the paper, we leave them out. Here we only mention that with these conventions, the Conley–Zehnder index of a non-degenerate minimum with small Hessian in dimension $2n$ is $n$.
The characteristic of the ground field $\mathbb{F}$} is essential in some parts of the paper. When the characteristic is zero, we can and usually will take $\mathbb{F}$$\mathbb{Q}$}. We use the notation ${\mathbb{F}}_{p}$ for the positive characteristic field ${\mathbb{Z}}/p{\mathbb{Z}}$, where $p$ is prime, or any other field of characteristic $p$.
@@ 2.2. Graded persistence modules
Recall furthermore that a (graded) interval persistence module ${\mathbb{F}}_{(a,\,b]}$, where $\infty<a<b\leq\infty$, is defined by setting
@@
The normal form or structure theorem asserts that every persistence module meeting the above conditions can be decomposed as a direct sum of a countable collection (i.e., a countable multiset) of interval persistence modules. Moreover, this decomposition is unique up to re-ordering of the sum. We refer the reader to [PRSZ] for a proof of this theorem and further references, and also, e.g., to [CZCG, ZC] for previous or related results. The multiset ${\mathcal{B}}(V)$ of intervals entering this decomposition is referred to as the barcode of $V$ and the intervals occurring in ${\mathcal{B}}(V)$ as bars. In addition, every bar is assigned a degree which is by definition the degree of the component of $V^{s}$ it comes from.
@@ 2.3. Filtered symplectic homology
The first versions of symplectic homology were introduced in [CFH, Vi99] and since then the theory has been further developed. We refer the reader to, e.g., [Abu, Bo, BO09a, BO09b, BO13, Se06] for the definitions and basic constructions, which we mainly omit or just informally sketch for the sake of brevity, focusing on lesser known facts. Here, treating equivariant and non-equivariant symplectic homology as persistence modules, we closely follow [ÇG2M25, GG20].
@@
This family of vector spaces together with the natural “inclusion” maps form a graded persistence module in the sense of Section 2.2 with ${\mathcal{S}}={\mathcal{S}}(\alpha)\cup\{0\}$. (It is essential at this point that $W$ meets the background requirement (BA). Otherwise, we would need to modify the definition of a persistence module; see [ÇG2M25].) We denote the resulting persistence module by $\operatorname{SH}(W;{\mathbb{F}})$. The global symplectic homology is then $\operatorname{SH}^{\infty}(W;{\mathbb{F}})$.
Somewhat informally, $\operatorname{SH}^{t}(W;{\mathbb{F}})$ can be defined via the following construction. As is pointed out in Section 2.2, it suffices to define $\operatorname{SH}^{t}(W;{\mathbb{F}})$ when $t\not\in{\mathcal{S}}$. Assume first that the Reeb flow is non-degenerate. Let $H$ be an autonomous Hamiltonian on $\widehat{W}$, which is a convex, strictly increasing function $h(r)$ of $r$ on $M\times[1,\infty)$ and 0 on $W$. The function $h$ is required to be linear at infinity, i.e., of the form trctr-c with $c>0$. Furthermore, let $\tilde{H}$ be a small Morse perturbation of $H$ supported near $W$. Then $\tilde{H}$ is a Morse–Bott Hamiltonian in the sense of [Bo, BO09a], and $\operatorname{SH}^{t}(W;{\mathbb{F}})$ is simply the Floer homology of $H$ or equivalently $\tilde{H}$; cf. [ÇG2M25, Thm. 3.5 and Cor. 3.6]. This is the homology of a complex with two generators $\hat{x}$ and $\check{x}$ of degrees, $\mu(x)+1$ and $\mu(x)$ respectively, for every closed Reeb of orbit $x$ with ${\mathcal{A}}(x)<t$ and one generator for every critical point of $\tilde{H}$; see, e.g., [Bo, BO09a, ÇG2M23]. In fact, one can trim down this complex further and make the number of generators of the latter type equal to $\dim\operatorname{H}(W;{\mathbb{F}})$. In any event, when $W$ is star-shaped we can have only one such generator, which we denote by $[W]$. When the Reeb flow is degenerate, we first take a small non-degenerate perturbation of the contact form and then apply this construction. For an interval $I=(t^{\prime},t)$ we define $\operatorname{SH}^{I}(W;\alpha)$ as the quotient of $\operatorname{SH}^{t}(W;{\mathbb{F}})$ by the image of $\operatorname{SH}^{t^{\prime}}(W;{\mathbb{F}})$.
@@
$\operatorname{SH}^{\delta}(W;{\mathbb{F}}):=\operatorname{H}(W;{\mathbb{F}})[-n];$ (2.2)
see, e.g., [BO09b, Vi99]. It readily follows from the above description of the symplectic homology as the Floer homology of $\tilde{H}$ that the Euler characteristic of the graded vector space $\operatorname{SH}^{t}(W;{\mathbb{F}})$ is independent of $t>0$. To be more precise, for any field $\mathbb{F}$} and every $t>0$, we have
$\sum_{m}(-1)^{m}\dim\operatorname{SH}^{t}_{m}(W;{\mathbb{F}})=(-1)^{n}\chi(W).$ (2.3)
Less directly, one can also use (2.11), (2.19) and (2.20) to prove (2.3).
The construction of symplectic homology carries over to $\mathbb{Z}$} or any other coefficient ring (see, e.g., [Abu]), and hence we have the universal coefficient theorem. In particular, for a fixed $t$, in every degree $m$ we have
@@
when the characteristic $p$ is large enough (depending on $t$). Indeed, the torsion part of the group $\operatorname{SH}^{t}_{m}(W;{\mathbb{Z}})$ is finite and (2.4) holds whenever $\operatorname{SH}^{t}_{m}(W;{\mathbb{Z}})$ has no elements of order $p$.
@@
Recall in this connection that $\operatorname{SH}^{\infty}(W;{\mathbb{F}})$ vanishes, for instance, whenever $W$ is displaceable in $\widehat{W}$. For example, $\operatorname{SH}(W)^{\infty}=0$ when $W$ is a star-shaped domain in $\widehat{W}={\mathbb{R}}^{2n}$. In this case, vanishing of $\operatorname{SH}^{\infty}(W)$ is established in [Vi99]. The general case is proved in [CFO] via Rabinowitz–Floer homology and direct proofs are given in [Su16] and [GS]. Strictly speaking, Theorem 2.3 is proved in [GS] for $\mathbb{F}$$\mathbb{Q}$}, but the argument carries over word-for-word to any unital ring.
@@
For every Liouville domain $W$ and any interval $I$ with end-points outside ${\mathcal{S}}(\alpha)$,
@@
These inequalities were originally proved for the Floer homology of any Hamiltonian $H$ on $\widehat{W}$ of the form H=trcH=tr-c outside a compact set with $t\not\in{\mathcal{S}}(\alpha)$, [Se15, ShZh]. The version stated here readily follows from those results combined with the fact that for any fixed interval $I$ with end-points outside ${\mathcal{S}}(\alpha)$ and a fixed prime $p$, we have $\operatorname{SH}^{I}(W;{\mathbb{F}}_{p})=\operatorname{HF}^{I}(H;{\mathbb{F}}_{p})$ for a suitable choice of $H$; see [ÇG2M23, Sect. 3]. For instance, for $I=(-\infty,t)$ we can take any $H$ such that H=trcH=tr-c at infinity. Then, $\operatorname{SH}^{I}(W;{\mathbb{F}}_{p})=\operatorname{HF}(H;{\mathbb{F}}_{p})$ by [ÇG2M25, Cor. 3.7], which implies (2.6).